OpenStudy (anonymous):

how many molecules are in each section of 4.0g H2O

OpenStudy (ciarán95):

To find out the number of molecules which are present in the 4.0g sample, we must first find out how many moles of H2O we have. From the periodic table, we can find the relative molecular mass of water. - The relative atomic mass of H is 1 g/mole (i.e: 1 mole of H atoms = 1 g). -The relative atomic mass of O (when rounded up) is 16 g/mole. -So, the relative molecular mass of H2O = 2(1 g/mole) + (16 g/mole) = 18 g/mole So, if we divide the amount of the sample we have (in grams) by the relative molecular mass, we should get the number of moles we have in the sample. So, you can work out 4.0 g/18 g/mole to find this value. We also know that 1 mole of a substance (element, compound, etc..) contains Avogadro's Number, or approximately 6.022 x 10^23 individual particles (atoms, molecules, etc...) of that substance. So, for example, if you had 0.5 moles of water (this is not correct for this question), and you wanted to find out how many individual molecules of water you had. then you could say: 1 mole = 6.022 x 10^23 molecules So, 0.5 moles = 0.5(6.022 x 10^23 molecules) = 3.011 x 10^23 molecules. If you follow these steps, then you should arrive at the answer. Hope that helps! :)